OpenStudy (anonymous):

Haji rows his canoe 10 mph in still water. He noticed one day that he could row 4 miles upstream in the same amount of time he could row 6 miles downstream. What was the approximate speed of the current that day?

OpenStudy (amistre64):

well, distance is equal to speed, times time. 4 = (10-c)t 6 = (10+c)t solving for t we would have: 4/(10-c) = 6/(10+c) solve for c

OpenStudy (amistre64):

the current reduces your rowing speed upstream, and adds to it going down stream ... hence the 10 +- c parts